# Benchmarking LLMs via Uncertainty Quantification

### v2

@@ 6.1 Setup
In our experiments, we set the error rate $\alpha$ to 0.1, implying that the prediction set should include the true label with a probability of at least 0.9. In order to better excite the ability of LLMs, we incorporate examples or demonstrations in the prompt, adhering to the in-context learning paradigm (Dong et al., 2022). Specifically, we provide five demonstrations for QA, RC, and CI tasks. For the DRS task, we utilize three demonstrations, while we use only one demonstration for the DS task due to constraints on input length and inference cost. The maximum input length for all tasks is set to 2048 tokens. In addition, for each task, we allocate 50% of the data as the calibration set and the remaining 50% as the test set. We report results on the test set. These results represent the average value obtained from the two conformal score functions, namely, LAC and APS, as well as the three prompting strategies. It is noteworthy that while both LAC and APS fulfill the coverage guarantee requirement, they may produce prediction sets of varying sizes. By taking the average value of these two score functions, we aim to mitigate the influence of different score functions on the evaluation of uncertainty, thereby ensuring a more rigorous and reliable assessment.
@@
Table 1: The CR (%) results of LLMs with sizes ranging from 6B to 14B. These results represent the mean values of LAC and APS. The "Avg." column denotes the average performance across the five tasks.
LLMs Acc (%) $\uparrow$ SS $\downarrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1) 2.80(2) 1.74(1) 2.02(2) 1.94(1) 2.37(3) 2.17(1)
Yi-6B 57.57(4) 85.99(2) 76.50(2) 58.72(4) 66.06(1) 68.97(2) 3.20(5) 1.92(4) 1.88(1) 2.85(6) 1.96(1) 2.36(2)
Gemma-7B 62.24(2) 85.29(3) 73.58(3) 66.79(2) 40.80(7) 65.74(3) 2.72(1) 1.88(3) 2.04(3) 2.14(2) 3.11(7) 2.38(3)
Mistral-7B 60.44(3) 81.94(5) 62.93(5) 53.21(5) 62.16(2) 64.14(4) 2.80(2) 1.75(2) 2.48(5) 2.71(5) 2.40(4) 2.43(4)
Llama-2-13B 52.52(6) 77.23(6) 59.66(6) 52.65(6) 60.05(3) 60.42(5) 3.06(4) 2.24(7) 2.72(6) 2.55(4) 2.24(2) 2.56(5)
Qwen-7B 55.21(5) 83.89(4) 63.70(4) 64.04(3) 32.53(9) 59.87(6) 3.26(7) 2.15(5) 2.28(4) 2.51(3) 2.92(5) 2.63(6)
InternLM-7B 48.37(7) 73.86(7) 46.21(7) 43.72(7) 34.38(8) 49.31(7) 3.49(9) 2.19(6) 3.28(9) 3.63(10) 4.47(11) 3.41(9)
Llama-2-7B 45.60(9) 65.79(8) 43.05(8) 32.61(9) 45.60(5) 46.53(8) 3.20(5) 2.39(8) 3.27(8) 3.26(7) 3.30(8) 3.09(7)
DeepSeek-7B 45.65(8) 65.39(9) 42.66(9) 33.50(8) 42.15(6) 45.87(9) 3.34(8) 2.77(9) 3.06(7) 3.40(8) 3.08(6) 3.13(8)
MPT-7B 29.49(10) 31.69(10) 25.50(10) 24.38(11) 24.86(10) 27.18(10) 3.53(10) 3.46(10) 3.60(10) 3.59(9) 3.66(9) 3.57(10)
Falcon-7B 23.75(11) 24.98(11) 24.91(11) 25.86(10) 24.69(11) 24.84(11) 3.90(11) 3.60(11) 3.66(11) 3.64(11) 3.92(10) 3.75(11)
Table 2: The Acc and SS results of LLMs with sizes ranging from 6B to 14B. These results represent the mean values of LAC and APS. The "Avg." column denotes the average performance across the five tasks. The small number in parentheses next to each score indicates the ranking of the LLM for that specific task. The relative ranking of LLMs is also visually demonstrated by the depth of color, with darker colors signifying higher rankings.
6.2 Evaluated Models
We select a diverse set of nine representative models (or model series) from the vast array of open-source LLMs available. These models encompass various architectures and training methodologies, thereby allowing for a comprehensive benchmarking analysis. Specifically, the chosen models include the Llama-2 series (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Falcon series [^5] We omit Falcon-180B due to insufficient GPU resources. (Almazrouei et al., 2023), MPT-7B (Team, 2023b), Gemma-7B (Team et al., 2024), Qwen series (Bai et al., 2023), Yi series (01.AI, 2023), DeepSeek series (DeepSeek, 2023), and InternLM-7B (Team, 2023a). For all models, we utilize their checkpoints from the HuggingFace platform [^6] https://huggingface.co/models.
6.3 Main Findings
In our primary experiments, we focus on LLMs with sizes ranging from 6B to 14B parameters. The outcomes of CR, Acc and SS are presented in Tables 1 and 2, respectively. As previously mentioned, the reported results are the mean value derived from the two conformal score functions, LAC and APS. For a detailed analysis of the results pertaining to each score function, please refer to Appendix C.1.
Figure 3: Performance comparison of different versions of the Qwen series (1.8B to 72B) on various tasks.
@@ 6.3 Main Findings
From Table 1, it is evident that in the majority of cases, the coverage rate is at least 90%, indicating that the coverage guarantee requirement has been met. Although there are cases where the coverage rate falls below 90% [^7] While the theoretical guarantee of conformal prediction is rigorous, there can be minor fluctuations in practice due to finite-sample variability (Angelopoulos and Bates, 2021)., the values are still in close proximity to the 90% threshold. The lowest coverage rate is attained by Qwen-7B on the DS task, with a value of 89.56%. Furthermore, all models achieve an average coverage rate exceeding 90% across the five tasks. These findings suggest that the generated prediction sets are meaningful, as they cover the true label with a high probability. Therefore, the size of the prediction set can serve as a reliable indicator of uncertainty.
@@
In principle, an LLM having higher accuracy should demonstrate lower uncertainty. However, as shown in Table 2, the results regarding the SS metric reveal that in practice, higher accuracy does not necessarily correlate with lower uncertainty. Concretely, for each task, we observe that the ranking of LLMs based on accuracy differs from that based on uncertainty, suggesting that some LLMs possessing higher accuracy actually display higher uncertainty. Notably, two LLMs with a significant difference in accuracy may even display inverse uncertainty. For example, on the DRS task, InternLM-7B demonstrates higher accuracy than MPT-7B in accuracy by 19.34 absolute points, yet it shows higher uncertainty. This pattern is also observed on the QA task with Qwen-7B and Llama-2-7B (9.61), the CI task with Qwen-14B and Yi-6B (14.50), and the DS task with InternLM-7B and Falcon-7B (9.69). In each case, the LLM with much higher accuracy exhibits greater uncertainty compared to its counterpart. These observations underscore the importance of considering uncertainty in addition to accuracy when evaluating LLMs.
@@ 6.5 Effects of Instruction Finetuning
LLMs CR (%) Acc (%) $\uparrow$ SS $\downarrow$
Mixtral-8x7B 92.59 73.74 2.03
Mistral-7B 91.78 64.14 2.43
Llama-2-13B 91.60 60.42 2.56
Qwen-14B 92.94 74.00 2.17
6.6 Effects of Mixture of Experts
A Mixture of Experts (MoE) is a technique that combines multiple specialized models with a gating mechanism to enhance performance by leveraging diverse expertise and adaptability in handling complex data relationships. In recent months, the adoption of the MoE technique to augment the performance of LLMs has been steadily increasing. Considering this, we study how MoE impacts the uncertainty of LLMs, specifically comparing the Mixtral-8x7B MoE model [^9] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 with Mistral-7B. The results are reported in Table 3. While Mixtral-8x7B has 46.7B total parameters, it only uses 12.9B parameters per token. It, therefore, processes input and generates output at the same speed and for the same cost as a 12.9B model [^10] https://mistral.ai/news/mixtral-of-experts/. To provide a comprehensive comparison, we also include the results of Llama-2-13B and Qwen-14B. As can be observed, Mixtral-8x7B consistently outperforms Mistral-7B across both accuracy and uncertainty. Remarkably, it achieves accuracy levels comparable to Qwen-14B, while demonstrating the lowest average uncertainty among the four LLMs. These observations show that MoE is indeed an effective method to enhance the accuracy and reduce the uncertainty of LLMs.
6.7 Effects of Amount of Calibration Data
As described in Section 3, conformal prediction requires a calibration set to calculate the threshold $\hat{q}$. In our prior analyses, we have allocated 50% of the data as the calibration set and the remaining 50% as the test set. Here, we explore the impact of varying the proportion of calibration data, ranging from 10% to 50%, on uncertainty quantification. Note that the same 50% of data is consistently used as the test set. The average results across five tasks of the Llama-2 series are shown in Figure 5. It is observed that there are no significant variations in coverage rate (CR) and uncertainty (SS) for all versions of the Llama-2 series when varying the amount of calibration data. This observation confirms the efficacy of applying conformal prediction for uncertainty quantification in our analysis.
LLMs CR (%) Acc (%) $\uparrow$ SS $\downarrow$
Yi-34B 94.22 81.46 1.93
Yi-34B 93.31 80.10 1.88
Llama-2-70B 93.11 72.24 2.16
Llama-2-70B 91.55 63.55 2.72
6.8 Ablation Study
Inspired by in-context learning (Xie et al., 2021), our previous analyses have incorporated demonstrations for each task. Here, we aim to compare the performance of LLMs when demonstrations are included or excluded from the prompt input. The results of Yi-34B and Llama-2-70B are reported in Table 4. We observe that the presence of demonstrations leads to improved performance for both models, as evidenced by increased accuracy. Having demonstrations also leads to lower uncertainty for Llama-2-70B but higher uncertainty for Yi-34B. Overall, these results substantiate the effectiveness of incorporating demonstrations into the prompt to stimulate the capabilities of LLMs.
6.9 Comparison to Other Uncertainty Quantification Methods
@@
Given the predicted probabilities, another widely used measure of uncertainty is entropy (Rényi, 1961). There have also been some entropy-based uncertainty quantification methods for language models (Fadeeva et al., 2023). Here, we compare conformal prediction to entropy. Since conformal prediction uses the prediction set size (SS) to measure uncertainty, a direct comparison with entropy is not straightforward. To address this issue, we convert entropy to perplexity (Jelinek et al., 1977), which is defined as $PPL=2^{H}$, where $H$ denotes the entropy. Perplexity takes values in the range of $[1,|\mathcal{Y}|]$, allowing it to be interpreted as prediction set size. For instance, when the predicted probability for each class (option) is $\frac{1}{|\mathcal{Y}|}$, $PPL=|\mathcal{Y}|$.
The results regarding InternLM-7B are presented in Table 5, from which, we observe that the coverage rate of perplexity varies significantly across different tasks. On the QA task, the coverage rate is only $83.44\%$. In contrast, the coverage rate of conformal prediction consistently exceeds $90\%$. These results demonstrate the advantages of using conformal prediction for uncertainty quantification. [^11] It is noted that except for coverage rate (CR) and prediction set size (SS), some other metrics such as AUROC (Kuhn et al., 2023) and PRR (Malinin and Gales, 2020) have been proposed to evaluate the performance of uncertainty quantification methods. However, it has been pointed out that these metrics have many limitations (Huang et al., 2024). Instead, both CR and SS are intuitive and highly interpretable, making them reasonable choices.
We provide more analyses in Appendix C.5.
@@ 7 Conclusion
In this work, we have provided an extensive examination of the performance of LLMs by focusing on prediction uncertainty. To achieve this, we have employed conformal prediction for uncertainty quantification. Our comprehensive investigation, which involves nine LLMs (or LLM series) and spans five tasks, demonstrates that relying solely on accuracy for benchmarking LLMs is insufficient. Instead, it is imperative to take uncertainty into account when assessing their overall performance.
Limitations
Despite the multiple advantages of conformal prediction over other uncertainty quantification methods, it demonstrates three key limitations when employed to assess the uncertainty of LLMs. Firstly, the application of conformal prediction necessitates access to model output logits, which precludes the possibility of benchmarking LLMs such as ChatGPT that are only accessible via their APIs. Secondly, the adoption of conformal prediction poses challenges in evaluating the generative capabilities of LLMs. In our analyses, all tasks are transformed into multiple-choice questions, thereby primarily assessing the language understanding abilities of LLMs rather than their generative potential. Thirdly, the prediction sets generated by conformal prediction could be significantly influenced by the conformal score function utilized. Thus, for a specific LLM, varying conformal score functions may yield disparate estimations of uncertainty.
Nevertheless, the limitations of other uncertainty quantification methods prevent us from applying them in the context of LLMs. Per our knowledge, conformal prediction is currently the most feasible technique for robust uncertainty quantification of LLMs. Furthermore, some recent studies have tried to incorporate conformal prediction into the language generation process (Ravfogel et al., 2023; Quach et al., 2023; Deutschmann et al., 2023). We posit that conformal prediction will eventually evolve into an appropriate method for quantifying the uncertainty of language generation in the future.
Last but not least, it is important to note that the scope of this study is limited to evaluating the capabilities of LLMs in the context of language processing exclusively. The current trend in the field is towards the development of multi-modal foundation models (Liu et al., 2023; Li et al., 2023b; Lyu et al., 2023), which have the capacity to process multiple modalities rather than just language. Therefore, it would be a significant extension of this research to investigate the uncertainty associated with these foundation models when they are applied to non-language modalities, which constitutes an important avenue for future research.
Ethics Statement
Quantifying uncertainties in LLMs is vital for ensuring their safe, reliable, and beneficial use in practical applications, such as informed decision-making. By incorporating uncertainty quantification through conformal prediction, this research contributes to a more nuanced understanding of LLMs and their applicability in various tasks. As a result, we hope this work can serve as a foundation for future research in the area of LLMs, emphasizing the importance of evaluating models from multiple perspectives to obtain a more comprehensive assessment of their performance.

### v3

@@ 6.1 Setup
In our experiments, we set the error rate $\alpha$ to 0.1, implying that the prediction set should include the true label with a probability of at least 0.9. In order to better excite the ability of LLMs, we incorporate examples or demonstrations in the prompt, adhering to the in-context learning paradigm [19]. Specifically, we provide five demonstrations for QA, RC, and CI tasks. For the DRS task, we utilize three demonstrations, while we use only one demonstration for the DS task due to constraints on input length and inference cost. The maximum input length for all tasks is set to 2048 tokens. In addition, for each task, we allocate 50% of the data as the calibration set and the remaining 50% as the test set. We report results on the test set. These results represent the average value obtained from the two conformal score functions, namely, LAC and APS, as well as the three prompting strategies. It is noteworthy that while both LAC and APS fulfill the coverage guarantee requirement, they may produce prediction sets of varying sizes. By taking the average value of these two score functions, we aim to mitigate the influence of different score functions on the evaluation of uncertainty, thereby ensuring a more rigorous and reliable assessment.
6.2 Evaluated models
We select a diverse set of nine representative models (or model series) from the vast array of open-source LLMs available. These models encompass various architectures and training methodologies, thereby allowing for a comprehensive benchmarking analysis. Specifically, the chosen models include the Llama-2 series [63], Mistral-7B [34], Falcon series [^3] We omit Falcon-180B due to insufficient GPU resources. [3], MPT-7B [62], Gemma-7B [60], Qwen series [7], Yi series [1], DeepSeek series [15], and InternLM-7B [61]. For all models, we utilize their checkpoints from the HuggingFace platform: https://huggingface.co/models.
6.3 Main findings
In our primary experiments, we focus on LLMs with sizes ranging from 6B to 14B parameters. The outcomes of CR, Acc and SS are presented in Table 6.3. As previously mentioned, the reported results are the mean value derived from the two conformal score functions, LAC and APS. For a detailed analysis of the results pertaining to each function, please refer to Appendix C.1.
Table 1: The evaluation results of LLMs with sizes ranging from 6B to 14B. These results represent the mean values of LAC and APS. The "Avg." column denotes the average performance across the five tasks. The small number in parentheses next to each score indicates the ranking of the LLM for that specific task. The relative ranking of LLMs is also visually demonstrated by the depth of color, with darker colors signifying higher rankings.
@@ 6.3 Main findings
Prediction Accuracy – Acc (%) $\uparrow$
[…]
Yi-6B 57.57(4) 85.99(2) 76.50(2) 58.72(4) 66.06(1) 68.97(2)
Gemma-7B 62.24(2) 85.29(3) 73.58(3) 66.79(2) 40.80(7) 65.74(3)
Mistral-7B 60.44(3) 81.94(5) 62.93(5) 53.21(5) 62.16(2) 64.14(4)
Llama-2-13B 52.52(6) 77.23(6) 59.66(6) 52.65(6) 60.05(3) 60.42(5)
Qwen-7B 55.21(5) 83.89(4) 63.70(4) 64.04(3) 32.53(9) 59.87(6)
InternLM-7B 48.37(7) 73.86(7) 46.21(7) 43.72(7) 34.38(8) 49.31(7)
Llama-2-7B 45.60(9) 65.79(8) 43.05(8) 32.61(9) 45.60(5) 46.53(8)
DeepSeek-7B 45.65(8) 65.39(9) 42.66(9) 33.50(8) 42.15(6) 45.87(9)
MPT-7B 29.49(10) 31.69(10) 25.50(10) 24.38(11) 24.86(10) 27.18(10)
Falcon-7B 23.75(11) 24.98(11) 24.91(11) 25.86(10) 24.69(11) 24.84(11)
Prediction Uncertainty – SS $\downarrow$
[…]
Yi-6B 3.20(5) 1.92(4) 1.88(1) 2.85(6) 1.96(1) 2.36(2)
Gemma-7B 2.72(1) 1.88(3) 2.04(3) 2.14(2) 3.11(7) 2.38(3)
Mistral-7B 2.80(2) 1.75(2) 2.48(5) 2.71(5) 2.40(4) 2.43(4)
Llama-2-13B 3.06(4) 2.24(7) 2.72(6) 2.55(4) 2.24(2) 2.56(5)
Qwen-7B 3.26(7) 2.15(5) 2.28(4) 2.51(3) 2.92(5) 2.63(6)
InternLM-7B 3.49(9) 2.19(6) 3.28(9) 3.63(10) 4.47(11) 3.41(9)
Llama-2-7B 3.20(5) 2.39(8) 3.27(8) 3.26(7) 3.30(8) 3.09(7)
DeepSeek-7B 3.34(8) 2.77(9) 3.06(7) 3.40(8) 3.08(6) 3.13(8)
MPT-7B 3.53(10) 3.46(10) 3.60(10) 3.59(9) 3.66(9) 3.57(10)
Falcon-7B 3.90(11) 3.60(11) 3.66(11) 3.64(11) 3.92(10) 3.75(11)
@@
From Table 6.3, it is evident that in the majority of cases, the coverage rate is at least 90%, indicating that the coverage guarantee requirement has been met. Although there are cases where the coverage rate falls below 90%, [^4] While the theoretical guarantee of conformal prediction is rigorous, there can be minor fluctuations in practice due to finite-sample variability [5]. the values are still in close proximity to the 90% threshold. The lowest coverage rate is attained by Qwen-7B on the DS task, with a value of 89.56%. Moreover, all models achieve an average coverage rate exceeding 90% across the five tasks. These findings suggest that the generated prediction sets are meaningful, as they can cover the true label with a high probability. Therefore, the size of the prediction set can serve as a reliable indicator of uncertainty.
@@
In principle, an LLM having higher accuracy is expected to demonstrate lower uncertainty. However, as shown in Table 6.3, the results regarding the SS metric reveal that in practice, higher accuracy does not necessarily correlate with lower uncertainty. Concretely, for each task, we observe that the ranking of LLMs based on accuracy differs from that based on uncertainty, suggesting that some LLMs possessing higher accuracy actually display higher uncertainty. Notably, two LLMs with a significant difference in accuracy may even display inverse uncertainty. For example, on the DRS task, InternLM-7B demonstrates higher performance than MPT-7B in accuracy by 19.34 absolute points, yet it shows higher uncertainty. This pattern is also observed on the QA task with Qwen-7B and Llama-2-7B (9.61), the CI task with Qwen-14B and Yi-6B (14.50), and the DS task with InternLM-7B and Falcon-7B (9.69). In each case, the LLM with much higher accuracy exhibits greater uncertainty compared to its counterpart. These observations underscore the importance of considering uncertainty in addition to accuracy when evaluating LLMs.
Figure 3: Performance comparison of different versions of the Qwen series (1.8B to 72B).
@@ 6.5 Effects of instruction finetuning
Table 3: Comparison among conformal prediction (CP), entropy (Entropy), and maximal predicted probability ($\mathbf{P_{max}}$) using InternLM-7B in terms of ECE (%).
Tasks CP Entropy $\mathbf{P_{max}}$
QA 15.83 15.83 15.83
RC 1.33 1.32 1.41
CI 3.16 3.45 3.75
DRS 12.11 12.40 12.45
DS 9.30 9.30 9.62
Avg. 8.35 8.46 8.61
6.6 Comparison to other uncertainty quantification methods
@@
Given the predicted probabilities, another widely used measure of uncertainty is entropy [56]. There have also been some entropy-based uncertainty quantification methods for language models [20]. Here, we compare conformal prediction to entropy. Since conformal prediction uses the prediction set size (SS) to measure uncertainty, a direct comparison with entropy is not straightforward. To address this issue, we convert entropy to perplexity [33], which is defined as $PPL=2^{H}$, where $H$ denotes the entropy. Perplexity takes values in the range of $[1,|\mathcal{Y}|]$, allowing it to be interpreted as prediction set size. For instance, when the predicted probability for each class (option) is $\frac{1}{|\mathcal{Y}|}$, $PPL=|\mathcal{Y}|$.
The results regarding InternLM-7B are presented in Table 3, from which, we observe that the coverage rate of perplexity varies significantly across different tasks. On the QA task, the coverage rate is only $83.44\%$. In contrast, the coverage rate of conformal prediction consistently exceeds $90\%$. This is because when measuring uncertainty, entropy doesn’t take accuracy into account. Entropy remains the same when predicted probabilities are permuted, even though prediction accuracy may differ.
To further demonstrate the superiority of conformal prediction, we conduct additional experiments comparing it with entropy and maximal predicted probability in terms of the Expected Calibration Error (ECE) metric [27]. The results corresponding to InternLM-7B are presented in Table 3. The observation that conformal prediction yields the lowest average ECE score suggests that it offers more reliable uncertainty quantification.
Overall, these results demonstrate the advantages of adopting conformal prediction for uncertainty quantification. We provide more analyses in Appendix C.7.
Table 4: The evaluation results of closed-source LLMs.
LLMs CR (%) Acc (%) SS
GPT-4 90.41 81.75 1.65
GPT-3.5 89.98 62.99 3.05
Qwen-72B (sampling) 90.54 70.29 2.43
Qwen-72B (logits) 93.34 73.55 2.33
6.7 Expanding benchmarking to closed-source LLMs
In this part, we extend our benchmarking from open-source LLMs to closed-source LLMs. While obtaining the exact output logits of closed-source LLMs is challenging, we can sample multiple answers and then estimate the probability of each choice. We perform an experiment on the MMLU (the QA task) dataset with GPT-3.5 and GPT-4 as the closed-source LLMs. To save cost, we only consider the base prompting strategy. Specifically, we first sample 50 answers for each question and calculate the frequency of each option. Then, we apply the softmax function with temperature scaling to prevent zero probabilities. To demonstrate the quality of this approximation, we also report the results of the open-source model Qwen-72B when getting its predictions via sampling and via logits, respectively. The results are shown in Table 4.
It is observed that GPT-4 demonstrates the highest accuracy and the lowest uncertainty. In addition, the average prediction set size (SS) of Qwen-72B (sampling) is relatively close to that of Qwen-72B (logits). For each question, we further calculate the Jensen-Shannon divergence (JSD) between the predictions of Qwen-72B (sampling) and Qwen-72B (logits). The average JSD is 0.05, indicating that the two predictions (estimated probability distributions) are highly similar. Therefore, we conclude that the approximation is of high quality.
Table 5: The evaluation results of free-form text generation on TriviaQA.
LLMs CR (%) Acc (%) SS
Qwen-72B 88.92 76.45 2.63
Llama-2-13B 83.89 71.83 2.40
Qwen-14B 82.79 66.57 3.83
Llama-2-7B 78.83 64.91 3.06
Qwen-7B 77.89 59.44 5.02
DeepSeek-7B 78.41 57.52 6.12
Falcon-7B 76.51 55.74 6.27
Table 6: The relationship between stratified prediction set size (SS) and prediction accuracy (Acc).
SS LAC APS Avg.
1 80.39 92.69 86.54
2 59.77 82.21 70.99
3 40.55 63.70 52.12
4 40.07 41.71 40.89
5 31.50 34.92 33.21
6 13.43 None 13.42
6.8 Expanding benchmarking to free-form text generation
Here, we further extend our benchmarking from multiple-choice question answering to free-form text generation. However, applying conformal prediction to text generation is a complex task due to the extensive range of potential responses. It is not feasible to compute the probability for each possible response and then use conformal prediction to select a subset. Nevertheless, many potential responses have a low probability of being generated, which allows us to reduce the selection space by sampling multiple generations.
Specifically, we adopt the TriviaQA dataset [35] (sampling 10,000 dev instances) for free-form text generation. We first generate 20 answers for each question. Then, we employ the perplexity [33] of each generation as the conformal score function and utilize exact match to verify the accuracy of the generated answer. The results are displayed in Table 6. Note that the value of SS falls into $[1,20]$.
From Table 6, we observe that the prediction set size (SS) varies among LLMs, which could provide some insights into the uncertainty of these models. However, we must note that in this sampling setting, the coverage rate cannot be guaranteed any more because there might not be a correct answer within the 20 sampled responses. In other words, even if the prediction set size is 20, indicating high model uncertainty, the coverage rate for that instance could still be zero if there are no correct answers present. Nonetheless, it is observed that when the LLM is stronger, the coverage guarantee requirement is more likely to be satisfied.
6.9 In-depth analysis of the set size metric in relation to prediction accuracy
While our main focus is on the high probability of the prediction set covering the ground truth, it is also insightful to explore the relationship between stratified set size and prediction accuracy. In our experiments, we employ InternLM-7B on the QA task, grouping instances by their predicted set size and reporting the accuracy within each group in Table 6. The results reveal that instances with smaller set sizes are generally associated with higher prediction accuracy, indicating that set size serves as a useful indicator of prediction uncertainty. However, it is important to note that even when the set size reaches its maximum value, there are still instances where the prediction is accurate. Consequently, a comprehensive analysis of both prediction accuracy and prediction uncertainty is essential for a thorough assessment of the performance of LLMs.
@@ 7 Conclusion
In this work, we have provided an extensive examination of the performance of LLMs by focusing on prediction uncertainty. To achieve this, we have employed conformal prediction for uncertainty quantification. Our comprehensive investigation, which involves nine open-source LLMs (or LLM series) and spans five typical NLP tasks, demonstrates that relying solely on accuracy for benchmarking LLMs is insufficient. Instead, it is imperative to take uncertainty into account when assessing their overall performance. Last but not least, we have verified the superiority of conformal prediction compared to several other uncertainty quantification methods. We have also extended our analyses to closed-source LLMs and free-form text generation.
Acknowledgments and Disclosure of Funding
This work was supported in part by the Tencent AI Lab Rhino-Bird (Grant No. EF2023-00151-FST), the Science and Technology Development Fund, Macau SAR (Grant Nos. FDCT/060/2022/AFJ, FDCT/0070/2022/AMJ), and the Multi-year Research Grant from the University of Macau (Grant No. MYRG-GRG2023-00006-FST-UMDF). We thank all reviewers for their precious comments.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The main claims made in the abstract and introduction are consistent with the results presented in § 6.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix E.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix F.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Our implementation is available at https://github.com/smartyfh/LLM-Uncertainty-Bench.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See § 6.1.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix A.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See § 4.
    2. (b)
      Did you mention the license of the assets? [Yes] See the supplementary materials.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We have made some modifications to the original datasets and also included our code.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No] All datasets used are publicly available.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]